OpenStudy (anonymous):

Factor completely p^4 - 625

OpenStudy (blockcolder):

\[\begin{align} p^2-625 &=(p^2)^2-(25)^2\\ &=(p^2+25)(p^2-25)\\ &=(p^2+25)(p+5)(p-5) \end{align}\]